Psychologist wants to determine if there is a linear relationship between the number of hours a person goes without sleep and the number of mistakes he/she makes on a simple test. The following data is recorded.

Hours without sleep 32 38 48 24 46 35 30 34 42
Number of Mistakes 6 8 13 5 7 6 5 8 12

Hours without sleep=x, Number of mistakes=y

(a) Find the equation of regression line. Round slope and y-intercept to nearest thousandth.  

(b) Find the residual when the number of hours is 32 hours.  Analyze the result.  

(c) Apply critical value test (not hypothesis testing).  Show inequality. By using critical value test, is there linear relation exists?  
Answer #1

(a) The equation of regression line is:

y = -2.7677 + 0.2885*x

(b) y = -2.7677 + 0.2885*32 = 6.464

Residual = 6 - 6.464 = -0.464

(c) The hypothesis being tested is:

H0: \rho = 0

Ha: \rho ≠ 0

t = 3.2

The critical value is 2.36.

Since 3.2 > 2.36, we can reject the null hypothesis.

Therefore, we can conclude that there is a linear relation.

0.594
r   0.771
Std. Error   1.979
n   9
k   1
Dep. Var. y
ANOVA table
Source SS   df   MS F p-value
Regression 40.1307 1   40.1307 10.24 .0151
Residual 27.4249 7   3.9178
Total 67.5556 8  
Regression output confidence interval
variables coefficients std. error    t (df=7) p-value 95% lower 95% upper
Intercept -2.7677
x 0.2885 0.0901 3.200 .0151 0.0753 0.5016